# Interval and $\ell$-interval rational parking functions

### v3

@@ 5.1. Preferential arrangements
Gross defines preferential arrangements in [Gross_preferential_arrangements] as the number of ways $n$ distinguishable objects can be distributed into $r$ distinguishable boxes where $r$ varies $1\leq r\leq n$. Now recall that Fubini rankings are rankings of $n$ competitors in which ties are allowed (Definition 2.3). There is a bijection between preferential arrangements and Fubini rankings, which we prove next in Theorem 5.2.
@@ 5.2. Unit interval rational parking functions and barred preferential arrangements
From [unit_pf], we know that the set of Fubini rankings are in bijection with the unit interval parking functions. In Theorem 5.2, we establish that Fubini rankings are also in bijection with preferential arrangements. Thus, preferential arrangements are in bijection with unit interval parking functions. In fact more is true. In this section, we establish that unit rational parking functions are in bijection with barred preferential arrangements.
Barred preferential arrangements were first introduced by Ahlbach et al. [ahlbachBarredPreferentialArrangements2013] and have been also studied by Nkonkobe and Murali [nkonkobe2019generalised] and they are a generalization of preferential arrangements, which we define next.
@@
  • Assume $i$ is the first value in $B_{j}$. Then $a_{i}=\mathrm{index}(\min(B_{j}))$. In this case car $i$ is the first car (and possibly the only car) with this preference. Implying that car $i$ parks in spot $\mathrm{index}(\min(B_{j}))$. This establishes that value $i$ appears in index $\mathrm{index}(\min(B_{j}))$ in $\mathcal{O}(\alpha)$.
@@
  • Assume $i$ is not the first value in $B_{j}$, but the $y$th value, with $y\geq 2$. Then $a_{i}=m_{j}+y-2=\mathrm{index}(\min(B_{j}))+y-2$. Note that by construction of $\alpha$, car $i$ parks in spot $\mathrm{index}(\min(B_{j}))+y-1$. Thus $i$ appears as entry $\mathrm{index}(\min(B_{j}))+y-1$ in $\mathcal{O}(\alpha)$. Now lets consider the index at which $i$ appears in $x(q)$. By assumption, $i$ appears as the $y$th entry in block $B_{j}$. This means that $i$ appears in $\mathrm{index}(\min(B_{j}))+y-1$ and, thus, $i$ appears as entry $\mathrm{index}(\min(B_{j}))+y-1$ in $x(q)$.
@@
We conclude with some enumerative consequences of Theorem 5.17 and the work of Ahlbach, Usatine, and Pippenger  [ahlbachBarredPreferentialArrangements2013].

### v4

@@ 5.1. Preferential arrangements
Gross defines preferential arrangements in [19] as the number of ways $n$ distinguishable objects can be distributed into $r$ distinguishable boxes where $r$ varies $1\leq r\leq n$. Now recall that Fubini rankings are rankings of $n$ competitors in which ties are allowed (Definition 2.3). There is a bijection between preferential arrangements and Fubini rankings, which we prove next in Theorem 5.2.
@@ 5.2. Unit interval rational parking functions and barred preferential arrangements
From [7], we know that the set of Fubini rankings are in bijection with the unit interval parking functions. In Theorem 5.2, we establish that Fubini rankings are also in bijection with preferential arrangements. Thus, preferential arrangements are in bijection with unit interval parking functions. In fact more is true. In this section, we establish that unit rational parking functions are in bijection with barred preferential arrangements.
Barred preferential arrangements were first introduced by Ahlbach et al. [3] and have been also studied by Nkonkobe and Murali [27] and they are a generalization of preferential arrangements, which we define next.
@@
  • Assume $i$ is the first value in $B_{j}$. Then $a_{i}=\mathrm{index}(\min(B_{j}))$. In this case car $i$ is the first car (and possibly the only car) with this preference. Implying that car $i$ parks in spot $\mathrm{index}(\min(B_{j}))$. This establishes that value $i$ appears in $\mathrm{index}(\min(B_{j}))$ in $\mathcal{O}(\alpha)$.
@@
  • Assume $i$ is not the first value in $B_{j}$, but the $y$th value, with $y\geq 2$. Then $a_{i}=m_{j}+y-2=\mathrm{index}(\min(B_{j}))+y-2$. Note that by construction of $\alpha$, car $i$ parks in spot $\mathrm{index}(\min(B_{j}))+y-1$. Thus $i$ appears as entry $\mathrm{index}(\min(B_{j}))+y-1$ in $\mathcal{O}(\alpha)$. Now lets consider the index at which $i$ appears in $x(q)$. By assumption, $i$ appears as the $y$th entry in block $B_{j}$. Thus, $i$ appears as entry $\mathrm{index}(\min(B_{j}))+y-1$ in $x(q)$.
@@
We conclude with some enumerative consequences of Theorem 5.17 and the work of Ahlbach, Usatine, and Pippenger  [3].